# Manipulating topology of quantum phase transitions by symmetry enhancement

### v1

@@ Abstract
Topology plays a pivotal role in the solid state: It provides an elegant understanding of phases and phase transitions beyond the Landau-Ginzburg-Wilson paradigm [1, 2, 3, 4, 5]. Specific examples that set the stage for this Letter include antiferromagnetic spin-$S$ chains, with low energy fluctuations described by an 1+1 dimensional SO(3) non-linear sigma model with topological $\theta$-terms at $\theta=2\pi S$. This provides a topological understanding of the distinct physics of half-integer and integer spin chains [2, 6]. By tuning the strength of the second nearest neighbor antiferromagnetic exchange coupling, the spin-1/2 chain can undergo a continuous phase transition to a valence bond solid (VBS) phase that breaks site-inversion symmetry [7, 8]. This transition is described by a level $k=1$ Wess-Zumino-Witten (WZW) theory with emergent SO(4) symmetry that locks in spin-spin and dimer-dimer correlations. The topological term accounts for the non-trivial structure of the VBS domain wall that hosts a spin-1/2 degree of freedom. The phase transition out of the VBS state occurs via the proliferation of domain walls and concomitant appearance of critical antiferromagnetic fluctuations. Importantly, the $k=1$ Wess-Zumino-Witten theory has a marginal SO(4) symmetry breaking operator the relevance of which depends upon its sign [9, 10]. Thereby the transition requires no fine tuning. Higher half-integer spin chains map onto $k=2S$ Wess-Zumino-Witten theories that correspond to multi-critical points such that fine tuning is necessary to study them [9, 11].
The generalization of the above to 2+1 dimensions has broadly lead to the theory of deconfined quantum criticality (DQC) [12, 13, 14, 15, 3, 16, 17]. The physics of the easy-plane DQC point (DQCP) as well as the criticality of the model presented in Ref. [18] can be understood in terms of a 2+1 dimensional SO(4) nonlinear sigma model with $\theta$-term at $\theta=\pi$ [16, 19]. On the other hand the DQCP, as realized in the $J-Q$ [20], loop [21, 22] or various fermionic models [23, 24], can map onto the 2+1 dimensional SO(5) nonlinear model with level $k=1$ WZW term [25, 26]. There is an ongoing discussion pertaining to the nature, weakly first order or continuous, of the aforementioned critical points [27, 28, 29]. For our purposes, the mere proximity of the models to a unitary critical point where topology, in terms of $\theta$-terms at $\theta=\pi$ or level $k=1$ WZW terms is of importance. The aim of this Letter is to propose a symmetry enhancement of the model of Ref. [23] so as to modify the topological aspects of the putative field theory to level $k=N$ WZW terms and $\theta$-terms at $\theta=N\pi$. Such a class of models, allows us to switch on and off the role of topology, and thereby between Landau-Ginzburg-Wilson and topological quantum phase transitions.
@@
Owing to time reversal symmetry, the action is real for $\lambda>0$ [39] such that $e^{-NS\left\{\bm{\Phi}(\bm{i},\tau)\right\}}$ can be used as a weight function for Monte Carlo importance sampling. Since the Hubbard-Stratonovich transformation does not break the SU($N$) flavour symmetry, simulations of the higher symmetry model merely require multiplying the action by $N$ [40]. As we will argue in the discussion section, the form of the action equally suggests that if the low energy physics at $N=1$ is dominated by a level $k=1$ WZW term then the corresponding critical point in the SU($N$) model is described by a level $k=N$ WZW term [25]. For all of the presented calculations we set $t=1$ and consider a half-filled band, i.e. one electron per site corresponding to a vanishing chemical potential. Moreover, we use a symmetric Trotter decomposition with an imaginary time discretization $\Delta\tau=0.1$ and employ a finite temperature code at $\beta=2L$. Looking ahead this choice is justified, since the symmetry breaking phases as well as the SM we encounter all have a dynamical exponent set by unity.
@@
Similar to the $N=1$ case we find a SM phase for small values of $\lambda$ as can be proposed out of Fig. 1(d) which shows the defining linear dependence of the uniform charge and spin susceptibility $\chi^{C/S}(\bm{q}=\bm{0})$ on temperature, approaching zero as temperature declines. As spin and charge are conserved in the Hamiltonian we calculate the uniform susceptibility by multiplying Eq. (4) by $\beta$ and summing over all the orbitals $\delta,\delta^{\prime}$. The QSH phase can be identified by considering the divergent behavior of the structure factor with respect to the system size as shown in Fig. 1(a). This can be supported by a vanishing $\chi^{C}$ indicating a charge gap alongside gapless spin excitations specified by a finite $\chi^{S}$.
The diverging structure factor of SSC correlations appearing for greater values of $\lambda$ viewed in Fig. 1(b) together with a finite $\chi^{C}$ reveal an SSC phase.
In contrast to the $N=1$ case, we observe the emergence of a Kekulé VBS phase that suppresses the SSC phase partially. This manifests again in the structure factor of Fig. 1(c) as well as a vanishing $\chi^{C}$ and $\chi^{S}$ indicating a trivial insulator. Additionally, the inset in Fig. 1(c) reveals the Kekulé pattern recovered by plotting the kinetic energy per bond upon introducing a pinning field to the hopping breaking translational symmetry. It can be obtained by calculating the expectation value of the VBS order parameter. Data for $\chi^{C/S}$ in the aforementioned ordered states are shown in the Supplemental Material [41].
@@
Figure 2: Correlation ratio of QSH order for the SM-QSH transition (a) together with a data collapse (b). Correlation ratio of QSH (c) and SSC (d) order parameters. The inset shows dynamical structure factors for $L=15$ at the QSH-SSC transition $\lambda=0.0975$. The slope denotes the respective Fermi velocity extracted from the Dirac cones in units of hopping amplitude $t$ times lattice constant $a$.
@@
In comparison to the $N=1$ phase diagram the model at $N=2$ shows an additional Kekulé VBS phase. It is interesting to view this result in the $N$-$\lambda$ plane. In the large-$N$ limit, owing to Eq. (3), one expects the saddle point approximation to be exact. As shown in the supplemental material [41] the SM phase at small $\lambda$ gives way to the QSH phase with mean-field order parameter exponent $\beta=1$. At finite values of $N$, one can account for the SSC phase by expanding the perfect square term, thereby revealing the presence of an attractive interaction in the s-wave channel [31]. Another way of understanding the SSC phase, is via the use of a Fierz identity (see Ref. [47] and Supplemental Material[41]). In particular since the QSH and SSC order parameters anti-commute the QSH mass term naturally also allows for an SSC instability. The Kekulé phase is surrounded by the SSC phases. The Kekulé and SSC mass terms are compatible (i.e. anticommute) such that one order potentially triggers the other. Importantly, the transitions we observe between ordered phases are between compatible (i.e. anticommuting) mass terms that in principle invalidate a Landau-Ginzburg-Wilson description: In $d+1$-dimensions, $d+3$ ($d+2$) compatible mass terms lead to WZW ($\theta$)- terms [19].
@@
We observe two transitions between the Kekulé and SSC phases. One is a first order transition, and the second shows coexistence. These are typical for Landau-Ginzburg-Wilson transitions between different symmetry broken phases, $U(1)$ charge and $U(1)$ rotational symmetry in the continuum. However, in our setting this is surprising since the two SSC, $M_{1}$, $M_{2}$ and two Kekulé, $M_{3}$, $M_{4}$ masses anticommute [50]. Let us start with the $N=1$ case, and supplement the Dirac vacuum with an interaction term $-U\sum_{i=1}^{4}(M_{i})^{2}$. Since $\Gamma_{i,j}=\frac{i}{2}\left[M_{i},M_{j}\right]$ correspond to the six generators of SO(4) the Hamiltonian possesses SO(4) global symmetry  [^1] A rotation $e^{i\theta\bm{e}_{\alpha}\Gamma_{\alpha}}\Psi(x)$ of the Dirac spinor, amounts to an SO(4) rotation of the vector of mass terms. . In the massive phase, one can integrate out the fermions to obtain a $\theta$-term at $\theta=\pi$ [19, 16] that accounts for the fact that the vortex of the U(1) Kekulé order parameter shows an emergent spin-1/2 degree of freedom. In Refs. [18, 52] such a transition with emergent SO(4) symmetry has been observed numerically and the consequences of the topological term were discussed in detail. We now have to account for the SU($N$) flavor symmetry. Importantly, this symmetry is not broken in our simulations (See supplemental material [41]). Under this condition, $N$ is merely a multiplicative factor of the action:
@@ Acknowledgements.
FFA would like to thank Lukas Janssen for discussions. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (www.lrz.de). FFA and TS thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490) and GR the DFG funded SFB 1170 on Topological and Correlated Electronics at Surfaces and Interfaces. MR acknowledges support by the DFG through Grant No. 332790403.

### v2

# Alteration of Topology in Quantum Phase Transitions via Symmetry Enrichment

@@ Abstract
Topology plays a pivotal role in the solid state: It provides an elegant understanding of phases and phase transitions beyond the Landau-Ginzburg-Wilson paradigm [1, 2, 3, 4, 5]. Specific examples that set the stage for this Letter include antiferromagnetic spin-$S$ chains, with low energy fluctuations described by an 1+1 dimensional SO(3) non-linear sigma model with topological $\theta$-terms at $\theta=2\pi S$. This provides a topological understanding of the distinct physics of half-integer and integer spin chains [2, 6]. By tuning the strength of the second nearest neighbor antiferromagnetic exchange coupling, the spin-1/2 chain can undergo a continuous phase transition to a valence bond solid (VBS) phase that breaks site-inversion symmetry [7, 8]. This transition is described by a level $k=1$ Wess-Zumino-Witten (WZW) theory with emergent SO(4) symmetry that locks in spin-spin and dimer-dimer correlations. The topological term accounts for the non-trivial structure of the VBS domain wall that hosts a spin-1/2 degree of freedom. The phase transition out of the VBS state occurs via the proliferation of domain walls and concomitant appearance of critical antiferromagnetic fluctuations. Importantly, the $k=1$ Wess-Zumino-Witten theory has a marginal SO(4) symmetry breaking operator the relevance of which depends upon its sign [9, 10]. Thereby the transition requires no multiparameter fine tuning. Higher half-integer spin chains map onto $k=2S$ Wess-Zumino-Witten theories that correspond to multi-critical points such that fine tuning is necessary to study them [9, 11].
The generalization of the above to 2+1 dimensions has broadly lead to the theory of deconfined quantum criticality (DQC) [12, 13, 14, 15, 3, 16, 17]. The physics of the easy-plane DQC point (DQCP) as well as the criticality of the model presented in Ref. [18] can be understood in terms of a 2+1 dimensional SO(4) nonlinear sigma model with $\theta$-term at $\theta=\pi$ [16, 19]. On the other hand the DQCP, as realized in the $J-Q$ [20], loop [21, 22] or various fermionic models [23, 24], can map onto the 2+1 dimensional SO(5) nonlinear model with level $k=1$ WZW term [25, 26]. There is an ongoing discussion pertaining to the nature, weakly first order or continuous, of the aforementioned critical points [27, 28, 29, 30]. For our purposes, the mere proximity of the models to a unitary critical point where topology, in terms of $\theta$-terms at $\theta=\pi$ or level $k=1$ WZW terms is of importance. The aim of this Letter is to propose a symmetry enhancement of the model of Ref. [23] so as to modify the topological aspects of the putative field theory to level $k=N$ WZW terms and $\theta$-terms at $\theta=N\pi$. Such a class of models, allows us to switch on and off the role of topology, and thereby between Landau-Ginzburg-Wilson and topological quantum phase transitions.
@@
Owing to time reversal symmetry, the action is real for $\lambda>0$ [40] such that $e^{-NS\left\{\bm{\Phi}(\bm{i},\tau)\right\}}$ can be used as a weight function for Monte Carlo importance sampling. Since the Hubbard-Stratonovich transformation does not break the SU($N$) flavour symmetry, simulations of the higher symmetry model merely require multiplying the action by $N$ [41]. As we will argue in the discussion section, the form of the action equally suggests that if the low energy physics at $N=1$ is dominated by a level $k=1$ WZW term then the corresponding critical point in the SU($N$) model is described by a level $k=N$ WZW term [25]. For all of the presented calculations we set $t=1$ and consider a half-filled band, i.e. one electron per site corresponding to a vanishing chemical potential. Moreover, we use a symmetric Trotter decomposition with an imaginary time discretization $\Delta\tau=0.1$ and employ a finite temperature code at $\beta=2L$. Looking ahead this choice is justified, since the symmetry breaking phases as well as the SM we encounter all have a dynamical exponent set by unity. The imaginary-time discretization $\Delta\tau$ leads to a systematic error that scales as $\Delta\tau^{2}$. This error must remain well below characteristic low-energy scales of the system to avoid contaminating physical observables. In the Supplemental Material of Ref. [23], it was shown that for $N=1$, choosing $\Delta\tau=0.2$ keeps such errors negligible. Meanwhile, Ref. [14] demonstrated that for larger $N$, setting $\Delta\tau\propto 1/N$ is a suitable guideline to ensure these errors remain below the relevant energy scale. Based on these results, we choose the aforementioned value of $\Delta\tau$ in our simulations.
@@
Similar to the $N=1$ case we find a SM phase for small values of $\lambda$ as can be proposed out of Fig. 1(d) which shows the defining linear dependence of the uniform charge and spin susceptibility $\chi^{C/S}(\bm{q}=\bm{0})$ on temperature, approaching zero as temperature declines. As spin and charge are conserved in the Hamiltonian the uniform susceptibility is given by $\chi^{O}(\bm{0})=\beta\sum_{\delta,\delta^{\prime}}S_{\delta,\delta^{\prime}}^{O}(\bm{0})$. The QSH phase can be identified by considering the divergent behavior of the structure factor with respect to the system size as shown in Fig. 1(a). This can be supported by a vanishing $\chi^{C}$ indicating a charge gap alongside gapless spin excitations specified by a finite $\chi^{S}$. Data for $\chi^{C/S}$ is shown in the Supplemental Material [42]. The diverging structure factor of SSC correlations appearing for greater values of $\lambda$ viewed in Fig. 1(b) together with a finite $\chi^{C}$ reveal an SSC phase.
In contrast to the $N=1$ case, we observe the emergence of a Kekulé VBS phase that suppresses the SSC phase partially. This manifests again in the structure factor of Fig. 1(c) as well as a vanishing $\chi^{C}$ and $\chi^{S}$ indicating a trivial insulator. Additionally, the inset in Fig. 1(c) reveals the Kekulé pattern recovered by plotting the kinetic energy per bond upon introducing a pinning field to the hopping breaking translational symmetry. It can be obtained by calculating the expectation value of the VBS order parameter.
@@
Figure 2: Correlation ratio of QSH order for the SM-QSH transition (a) together with a data collapse (b) employed to extract the values of $\lambda_{c}$ and $1/\nu$ by fitting to $L^{1/\nu}(\lambda-\lambda_{c})$. Correlation ratio of QSH (c) and SSC (d) order parameters. The inset shows dynamical structure factors for $L=15$ at the QSH-SSC transition $\lambda=0.0975$. The slope denotes the respective Fermi velocity extracted from the Dirac cones in units of hopping amplitude $t$ times lattice constant $a$.
@@
In comparison to the $N=1$ phase diagram the model at $N=2$ shows an additional Kekulé VBS phase. It is interesting to view this result in the $N$-$\lambda$ plane. In the large-$N$ limit, owing to Eq. (3), one expects the saddle point approximation to be exact. As shown in the supplemental material [42] the SM phase at small $\lambda$ gives way to the QSH phase with mean-field order parameter exponent $\beta=1$. At finite values of $N$, one can account for the SSC phase by expanding the perfect square term, thereby revealing the presence of an attractive interaction in the s-wave channel [32]. Another way of understanding the SSC phase, is via the use of a Fierz identity (see Ref. [48] and Supplemental Material [42]). In particular since the QSH and SSC order parameters anti-commute the QSH mass term naturally also allows for an SSC instability. The Kekulé phase is surrounded by the SSC phases. The Kekulé and SSC mass terms are compatible (i.e. anticommute) such that one order potentially triggers the other according to Fierz identity. In the Supplemental Material [42] we describe, that unlike other VBS phases observed in the literature following Ref. [49] the VBS observed here has no mean-field counterpart in the model. Here, the VBS vanishes together with the SSC as $N\to\infty$. Importantly, the transitions we observe between ordered phases are between compatible (i.e. anticommuting) mass terms that in principle invalidate a Landau-Ginzburg-Wilson description: In $d+1$-dimensions, $d+3$ ($d+2$) compatible mass terms lead to WZW ($\theta$)- terms [19].
@@
We observe two transitions between the Kekulé and SSC phases. One is a first order transition, and the second shows a coexistence regime which can be argued to be called a supersolid [52, 53, 54, 55, 56, 57, 58, 59, 60], whose thermodynamic stability can be confirmed by a finite uniform charge susceptibility as shown in the Supplemental Material [42]. These are typical for Landau-Ginzburg-Wilson transitions between different symmetry broken phases, $U(1)$ charge and $U(1)$ rotational symmetry in the continuum. However, in our setting this is surprising since the two SSC, $M_{1}$, $M_{2}$ and two Kekulé, $M_{3}$, $M_{4}$ masses anticommute [61]. Let us start with the $N=1$ case, and supplement the Dirac vacuum with an interaction term $-U\sum_{i=1}^{4}(M_{i})^{2}$. Since $\Gamma_{i,j}=\frac{i}{2}\left[M_{i},M_{j}\right]$ correspond to the six generators of SO(4) the Hamiltonian possesses SO(4) global symmetry  [^1] A rotation $e^{i\theta\bm{e}_{\alpha}\Gamma_{\alpha}}\Psi(x)$ of the Dirac spinor, amounts to an SO(4) rotation of the vector of mass terms. . In the massive phase, one can integrate out the fermions to obtain a $\theta$-term at $\theta=\pi$ [19, 16] that accounts for the fact that the vortex of the U(1) Kekulé order parameter shows an emergent spin-1/2 degree of freedom. In Refs. [18, 63] such a transition with emergent SO(4) symmetry has been observed numerically and the consequences of the topological term were discussed in detail. We now have to account for the SU($N$) flavor symmetry. Importantly, this symmetry is not broken in our simulations (See supplemental material [42]). Under this condition, $N$ is merely a multiplicative factor of the action:
@@
The processed data used for generating the figures of this Letter can be accessed under (DOI) 10.58160/t2n8a0h6n21k1mha and the code is available online [71].
@@ Acknowledgements.
FFA would like to thank Lukas Janssen for discussions. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre (www.lrz.de). Furthermore, the authors acknowledge HPC recources by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b133ae. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.. FFA and TS thank the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter ct.qmat (EXC 2147, project-id 390858490) and GR the DFG funded SFB 1170 on Topological and Correlated Electronics at Surfaces and Interfaces. MR acknowledges support by the DFG through Grant No. 332790403.